# Discovering Message Passing Hierarchies for Mesh-Based Physics Simulation

### v1

@@ Abstract
Graph neural networks have emerged as a powerful tool for large-scale mesh-based physics simulation. Existing approaches primarily employ hierarchical, multi-scale message passing to capture long-range dependencies within the graph. However, these graph hierarchies are typically fixed and manually designed, which do not adapt to the evolving dynamics present in complex physical systems. In this paper, we introduce a novel neural network named DHMP, which learns Dynamic Hierarchies for Message Passing networks through a differentiable node selection method. The key component is the anisotropic message passing mechanism, which operates at both intra-level and inter-level interactions. Unlike existing methods, it first supports directionally non-uniform aggregation of dynamic features between adjacent nodes within each graph hierarchy. Second, it determines node selection probabilities for the next hierarchy according to different physical contexts, thereby creating more flexible message shortcuts for learning remote node relations. Our experiments demonstrate the effectiveness of DHMP, achieving $22.7\%$ improvement on average compared to recent fixed-hierarchy message passing networks across five classic physics simulation datasets.
@@ 1 Introduction
Simulating physical systems with deep neural networks has achieved remarkable success due to their efficiency compared with traditional numerical solvers. Graph Neural Networks (GNNs) have been validated as a powerful tool for mesh-based physical scenarios, such as fluids and rigid collisions (Wu et al., 2020). The primary mechanism driving the GNN-based models is message passing (Sanchez-Gonzalez et al., 2020; Pfaff et al., 2021; Allen et al., 2023). In this process, time-varying physical quantities are encoded within the mesh structure at each time step, enabling each node to update its features by aggregating information broadcast from neighboring nodes. These existing methods generally rely on local message passing, which limits their ability to propagate influence over long distances. A common solution involves using multi-scale graph structures to facilitate direct information shortcuts between distant nodes (Lino et al., 2022; Cao et al., 2023; Yu et al., 2024; Han et al., 2022; Fortunato et al., 2022).
However, as illustrated in Table 1, these approaches typically depend on heuristic methods to create coarser message passing structures using predefined graphs (Cao et al., 2023; Yu et al., 2024) or downsample the nodes based on spatial proximity (Lino et al., 2022). These hierarchies are preprocessed in one pass before training. This results in fixed graph hierarchies over the entire physical sequence that do not account for the diverse range of physical contexts; while in practical systems like turbulence, despite identical boundary conditions, even minor changes in initial conditions can lead to significant differences in subsequent dynamics. Moreover, the spatial correlations in a physical process can evolve over time, making static GNN hierarchies insufficient for accommodating the time-varying node interactions.
To tackle the above challenge, we propose a novel deep learning approach named Dynamic Hierarchical Message Passing (DHMP), which constructs context-aware and temporally evolving graph hierarchies based on the original mesh topology and the input physical quantities. The key insight is to develop a differentiable node selection method that allows for flexible modeling of node interactions. This is technically supported by a straightforward yet effective anisotropic message passing strategy, which aggregates the neighboring features to the central node in a directionally non-uniform manner, predicting its downsampling probabilities as a node within the coarser graph level. We then employ Gumbel-Softmax sampling to create a differentiable approximation of the hard sampling process for the downsampled graph.
The anisotropic message passing mechanism not only adaptively creates multi-scale graph structures but also enables learned directionally non-uniform importance weights to facilitate both intra-level and inter-level propagation of dynamic information. As shown in Table 1, existing approaches perform isotropic feature aggregation within each message passing layer, assuming equal contributions from neighboring nodes, which may overlook the directional nature of physical processes. While some methods employ attention mechanisms to replace isotropic intra-level propagation (Dwivedi & Bresson, 2020; Janny et al., 2023; Yu et al., 2024; Han et al., 2022), our approach demonstrates advantages in computational efficiency. Furthermore, existing models generally rely on either predefined or unlearnable importance weights to transfer information across hierarchical levels. In contrast, the inter-level aggregation weights in DHMP are data-specific and time-varying, effectively harnessing the anisotropic nature of our message passing mechanism to enhance multi-scale modeling flexibility.
We evaluate our model on publicly available datasets, including CylinderFlow, Airfoil, Flag, and DeformingPlate, achieving $22.7\%$ improvement on average compared to recent fixed-hierarchy message passing networks. Additionally, DHMP demonstrates superior performance on challenging physical systems with time-varing mesh structures and unseen mesh resolutions.
Table 1: Comparison of mesh-based simulation models. Dynamic hierarchy refers to hierarchical graph structures evolving over time. Context-aware indicates that the graph structures are determined by the physical inputs. Prop. denotes different feature propagation mechanisms.
Model Hierachical Dynamic Context-Aware Anisotropic Learnable
Hierarchy Hierarchy Intra-level Prop. Inter-level Prop.
MGN (2020) - - -
BSMS-GNN (2023)
@@
DHMP
@@ 2 Preliminaries
where $\mathbf{v}_{i}$ is the feature of node $v_{i}\in\mathcal{V}$ and $\psi$ denotes a non-parametric aggregation function. The function $\phi^{e}$ updates the features of edges based on the endpoints, while $\phi^{v}$ updates the node states with aggregated messages from its neighbors. In existing mesh-based simulation methods, multi-layer perceptrons (MLPs) with residual connections are commonly employed for $\phi^{e}(\cdot)$ and $\phi^{v}(\cdot)$, with the aggregation function $\psi(\cdot)$ being defined as the sum of edge features. Notably, since the aggregation function treats all neighbors equally, the contributions from neighboring nodes may be averaged out, and the repeated message-passing process can further dilute distinctive node features. This issue is exacerbated in dynamic physical systems, where transferring directed patterns is crucial.
@@
To facilitate long-range modeling, hierarchical MPNNs process information at $L$ scales by creating a graph for each level and propagating information between them (Lino et al., 2022; Fortunato et al., 2022; Cao et al., 2023; Yu et al., 2024). Let $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ represent the graph structure at the finest level, i.e., the input mesh. The lower-resolution graphs $\mathcal{G}_{2},\mathcal{G}_{3},\ldots,\mathcal{G}_{L}$, with $|\mathcal{V}_{1}|>|\mathcal{V}_{2}|>\ldots>|\mathcal{V}_{L}|$, contain fewer nodes and edges, which allows for more efficient feature propagation over longer distances. The typical process for constructing multi-scale structures primarily involves downsampling and upsampling between adjacent graph hierarchies. Downsampling reduces the number of nodes while upsampling transfers information from a lower-resolution graph to a higher-resolution one. The downsampling operation can be broken down into two steps:
  • SELECT: Nodes are selected from the current graph structure $\mathcal{G}_{l}$ to create a new, coarser graph $\mathcal{G}_{l+1}$. Various criteria for node selection (Diehl, 2019; Ying et al., 2018; Lino et al., 2022) can be applied to form $\mathcal{V}_{l+1}$. The edges $\mathcal{E}_{l+1}$ in $\mathcal{G}_{l+1}$ are constructed by connecting the selected nodes based on the original edges $\mathcal{E}_{l}$. However, this process can sometimes may lead to loss of connectivity and introduce partitions Gao & Ji (2019); Lee et al. (2019); Cao et al. (2023). To mitigate this, connectivity in $\mathcal{E}_{l+1}$ can be strengthened by adding $K$-hop edges.
@@
The upsampling process is represented by the EXPAND, the inverse of the REDUCE function, which aggregates information from the coarser level back to the finer level. Most previous work generates coarser graphs for each sequence either by using numerical software or by downsampling the input mesh through heuristic pooling strategies. This process occurs during the data preprocessing stage, enabling the preprocessed hierarchy of the same input mesh topology to be reused across various initial conditions and different time steps.
@@ 3 Method
In this section, we introduce the Dynamic Hierarchical Message Passing Networks (DHMP), a fully differentiable model that learns to dynamically generate coarser graphs over the sequence while simultaneously learning to simulate the physical system over the learned hierarchical graphs.
3.1 Overview
Figure 1 demonstrates an overview of the proposed model, which operates in an encode-process-decode pipeline. The encoder first maps the input field to a latent feature space $\mathbf{V}_{1}=\{\mathbf{v}_{i}|v_{i}\in\mathcal{V}_{1}\}$ at the original mesh resolution. Subsequently, we model the physical dynamics across the learned multi-scale graph hierarchies with adaptive graph structures. To enhance the propagation of long-term dependencies between distant nodes, we propose an anisotropic message passing (AMP) mechanism, which is largely inspired by the directed nature of significant dynamic patterns.
In Section 3.2, we present the details of the AMP layer. In Section 3.3, we discuss the approach for learning context-aware graph hierarchies. In Section 3.4, we describe the inter-level downsampling and upsampling processes that incorporate AMP-based feature propagation. Finally, in Section 3.5, we outline the implementation details and hyperparameter choices.
@@ 3.2 Anisotropic Message Passing
We introduce the AMP layer, which facilitates information propagation both within and between graph hierarchies, enabling DHMP to effectively capture local and long-range dependencies simultaneously.
As shown in Eq. 1, a common method in mesh-based simulation is to use the summation aggregation function for node update: $\mathbf{\hat{v}}_{i}=\phi^{v}\Big{(}\mathbf{v}_{i},\sum_{v_{j}\in\mathcal{N}_{v_{i}}}\mathbf{\hat{e}}_{ij}\Big{)}$, where $v_{j}\in\mathcal{N}_{v_{i}}$ is a neighboring node of $v_{i}$ in the graph if $e_{ij}\in\mathcal{E}$. Using the summation aggregation has two drawbacks: i) it can excessively smooth the neighboring features, potentially failing to capture intricate local relations, as discussed in previous literature (Alon & Yahav, 2021; Dong et al., 2023; Dwivedi et al., 2022), and ii) it does not account for the directed nature that can be inherent in physics scenarios.
To address these issues, we propose the AMP layer, which employs a more flexible aggregation function to facilitate anisotropic feature propagation within each message-passing hierarchy. Instead of directly summing the edge features, AMP exploits learnable parameters $\phi^{w}:\mathbb{R}^{F^{e}}\rightarrow\mathbb{R}$ to predict the importance weight of edge feature $\mathbf{\hat{e}}_{ij}$ to node $v_{i}$:
$w_{ij}=\phi^{w}(\mathbf{e}_{ij},\mathbf{v}_{i},\mathbf{v}_{j}).$ (2)
To ensure that the coefficients are easily comparable across different nodes, we normalize them using the softmax function across all choices of $j$:
$\alpha_{ij}=\operatorname{softmax}_{j}\left(w_{ij}\right)=\frac{\exp\left(w_{ij}\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(w_{ik}\right)}.$ (3)
@@
Unlike traditional MPNNs with non-parametrized aggregation functions, the proposed AMP layer allows for the implicit assignment of varying contribution weights to the updated edge features within the same neighborhood. Analyzing the learned direction-specific weights in AMP provides additional benefits for interpretability. Besides, AMP differs from Graph Convolutional Networks (GCNs) (Niepert et al., 2016) and attention-based GNNs (Han et al., 2022; Janny et al., 2023) in its architectural design and the inclusion of relative node distances into edge features, which improves generalization and is crucial for capturing spatial relations. In contrast, GCNs and attention mechanisms primarily focus on node interactions based on absolute spatial positions.
Figure 1: In DHMP, physical dynamics is modeled on multiple graph resolutions with adaptive structures, $\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{L}$, and are processed using their respective AMP layers. The DiffSELECT operation performs differentiable pooling to create coarser graphs with learnable downsampling probabilities. REDUCE and EXPAND integrate inter-level information using learned feature aggregation weights over the neighboring nodes. DHMP is trained end-to-end with one-step supervision.
@@ 3.3 Differentiable Multi-Scale Graph Construction
In the DiffSELECT operation, we train the node update module $\phi^{v}$ based on anisotropic aggregated edge features to produce a probability $p_{i}$ for each node. This probability indicates the likelihood of retaining node $v_{i}$ in the next-level coarser graph $\mathcal{G}_{l+1}$. Accordingly, we rewrite Eq. 4 as follows:
$\mathbf{\hat{v}}_{i}^{l},p^{l}_{i}=\phi^{v}\Big{(}\mathbf{v}^{l}_{i},\sum_{v_{j}\in\mathcal{N}_{v_{i}}}\alpha^{l}_{ij}\mathbf{\hat{e}}^{l}_{ij}\Big{)}.$ (5)
Next, we employ Gumbel-Softmax sampling (Jang et al., 2017) on $p_{i}$ to determine whether node $v_{i}$ is included in the downsampled graph:
$z^{l}_{i}=\text{Gumbel-Softmax}\big{(}p^{l}_{i}\big{)},$ (6)
where $z^{l}_{i}$ is a binary variable indicating the selection of node $v_{i}$. When $z^{l}_{i}=1$, node $v_{i}$ is retained in the next graph level. In this way, the node set $\mathcal{V}_{l+1}$ is dynamically constructed based on node features from the finer graph level. The Gumbel-Softmax technique provides a differentiable approximation to hard sampling, thereby facilitating end-to-end training. Additionally, we implement the Gumbel-Softmax with temperature annealing to stabilize training, initially encouraging the exploration of hierarchies and gradually refining the selection process.
The edges $\mathcal{E}_{l+1}$ in the coarser graph $\mathcal{G}_{l+1}$ are constructed by connecting the selected nodes using the original graph’s edges $\mathcal{E}_{l}$. However, this process may result in disconnected partitions (see Figure 7 in the appendix). To address this issue, we enhance the connectivity in $\mathcal{E}_{l+1}$ by incorporating the $K$-hop edges during the edge selection process, defined as follows:

### v2

# EvoMesh: Adaptive Physical Simulation with Hierarchical Graph Evolutions

@@ Abstract
Graph neural networks have been a powerful tool for mesh-based physical simulation. To efficiently model large-scale systems, existing methods mainly employ hierarchical graph structures to capture multi-scale node relations. However, these graph hierarchies are typically manually designed and fixed, limiting their ability to adapt to the evolving dynamics of complex physical systems. We propose EvoMesh, a fully differentiable framework that jointly learns graph hierarchies and physical dynamics, adaptively guided by physical inputs. EvoMesh introduces anisotropic message passing, which enables direction-specific aggregation of dynamic features between nodes within each hierarchy, while simultaneously learning node selection probabilities for the next hierarchical level based on physical context. This design creates more flexible message shortcuts and enhances the model’s capacity to capture long-range dependencies. Extensive experiments on five benchmark physical simulation datasets show that EvoMesh outperforms recent fixed-hierarchy message passing networks by large margins. Code is available at https://github.com/hbell99/EvoMesh.
Machine Learning, ICML
@@ 1 Introduction
Table 1: Comparison of mesh-based physical simulation models. Dynamic hierarchy refers to hierarchical graph structures that evolve over time. Adaptive indicates that the graph structures are determined by physical inputs. Prop. denotes feature propagation.
Model Dynamic Adaptive Anisotropic Learnable
Hierarchy Hierarchy Intra-level Prop Inter-level Prop
MGN (2021)
@@
BSMS (2023)
Eagle (2023)
HCMT (2024)
EvoMesh
Simulating physical systems with deep neural networks has achieved remarkable success due to their efficiency compared with traditional numerical solvers. Graph Neural Networks (GNNs) have been validated as a powerful tool for mesh-based simulation, such as for fluids and rigid collisions (Wu et al., 2020). The primary mechanism driving the GNN-based models is message passing, where time-varying physical quantities are encoded within the mesh structure and are temporally updated by aggregating information broadcast from neighboring nodes (Sanchez-Gonzalez et al., 2020; Pfaff et al., 2021; Allen et al., 2023). Existing methods generally rely on repeated local message passing to propagate influence over long distances, which becomes extremely costly for large-scale mesh graphs. A common solution involves using multi-scale graph structures to create direct information shortcuts between distant nodes. (Lino et al., 2022; Cao et al., 2023; Yu et al., 2024; Han et al., 2022; Fortunato et al., 2022).
However, as shown in Table 1, previous methods commonly rely on heuristic node selection to create predefined (data-independent) coarser message passing graphs (Cao et al., 2023; Yu et al., 2024). These predefined graphs limit the model’s adaptation ability in two key ways. First, the fixed graph hierarchies, applied to the entire input sequence, do not account for the variety of physical contexts. In practical systems like turbulence, even with identical boundary conditions, small changes in initial conditions can lead to significant differences in subsequent dynamics. Second, since the spatial correlations in a physical process can evolve over time, static graph hierarchies are insufficient for capturing the time-varying node interactions.
To tackle this challenge, we propose a novel neural network approach named EvoMesh, which constructs data-adaptive and time-evolving graph hierarchies based on the input physical quantities. The key insight is to develop a differentiable node selection method that allows for flexible correlation of long-range, dynamic node interactions. This is technically supported by an anisotropic message passing (AMP) mechanism, which (i) aggregates neighboring features with non-uniform, learnable importance weights within each hierarchical level, (ii) predicts the probabilities of the node being retained for the next hierarchy based on physical context, and (iii) adaptively learns cross-hierarchy interactions to optimize information flow across scales. To enable differentiability in the node selection process, we approximate the discrete downsampling decisions using Gumbel-Softmax.
Another advantage of the AMP mechanism is its ability to enable features to transfer between nodes with varying importance, aligning with the directionally non-uniform nature of the dynamic patterns, as observed in scenarios such as CylinderFlow, AirFoil, and Flying Flag simulations. It applies to both intra-level and inter-level feature propagation. In contrast, as shown in Table 1, most previous GNN-based mesh simulation methods perform isotropic feature aggregation within the intra-level transition and rely on unlearnable importance weights to transfer inter-level information across hierarchical levels, assuming equal contributions from neighboring nodes.
Overall, our contributions are summarized as follows:
  • We present EvoMesh, which generates dynamic graph hierarchies through differentiable node selection, enabling adaptive modeling of multi-scale physical relations.
  • EvoMesh employs anisotropic message passing to enable directionally varied feature propagation both within and across graph hierarchies.
  • On average, EvoMesh outperforms fixed-hierarchy models by around $20\%$ across a range of standard benchmarks. It also demonstrates strong generalization to test cases with time-varying mesh structures, novel resolutions, and out-of-distribution dynamics.
@@ 2 Preliminaries
[…]
where $\mathbf{v}_{i}$ is the feature of node $v_{i}\in\mathcal{V}$ and $\psi$ denotes a non-parmatric aggregation function. The function $\phi^{e}$ updates the features of edges based on the endpoints, while $\phi^{v}$ updates the node states with aggregated messages from its neighbors. In existing GNN-based mesh simulation methods, multi-layer perceptrons (MLPs) with residual connections are commonly employed for $\phi^{e}(\cdot)$ and $\phi^{v}(\cdot)$, with the non-parametric aggregation function $\psi(\cdot)$ being defined as the sum of edge features. Notably, since the aggregation function treats all neighbors equally, the contributions from neighboring nodes may be averaged out, and the repeated message-passing process can further dilute distinctive node features. This issue is exacerbated in dynamic physical systems, where transferring directed patterns is crucial. Attention-based methods address this issue by reweighting neighbor features, either locally or globally (Veličković et al., 2018; Yu et al., 2024; Han et al., 2022; Yun et al., 2019). While effective for directional aggregation, most weighting remains limited to intra-level features and does not support dynamic graph hierarchy construction.
@@
To facilitate long-range modeling, hierarchical MPNNs process information at $L$ scales by creating a graph for each level and propagating information between them (Lino et al., 2022; Fortunato et al., 2022; Cao et al., 2023; Yu et al., 2024). Let $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ represent the graph structure at the finest level, i.e., the input mesh. The lower-resolution graphs $\mathcal{G}_{2},\mathcal{G}_{3},\ldots,\mathcal{G}_{L}$, with $|\mathcal{V}_{1}|>|\mathcal{V}_{2}|>\ldots>|\mathcal{V}_{L}|$, contain fewer nodes and edges, which allows for more efficient feature propagation over longer physical distances with certain propagation steps. The typical process for constructing multi-scale structures primarily involves downsampling and upsampling between adjacent graph hierarchies. Downsampling reduces the number of nodes while upsampling transfers information from a lower-resolution graph to a higher-resolution one. The downsampling operation includes two steps:
  • SELECT: Nodes are selected from the current graph structure $\mathcal{G}_{l}$ to create a new, coarser graph $\mathcal{G}_{l+1}$. Various strategies have been proposed to construct $\mathcal{V}_{l+1}$, including hand-crafted designs (Lino et al., 2022; Cao et al., 2023; Yu et al., 2024), geometric clustering (Han et al., 2022; Janny et al., 2023), and differentiable pooling methods that predict cluster assignments or select top-ranked informative nodes (Ying et al., 2018; Gao & Ji, 2019; Lee et al., 2019; Ranjan et al., 2020). The edges $\mathcal{E}_{l+1}$ in $\mathcal{G}_{l+1}$ are constructed by connecting the selected nodes based on the original edges $\mathcal{E}_{l}$. However, this process can sometimes lead to loss of connectivity and introduce partitions (Gao & Ji, 2019; Lee et al., 2019; Cao et al., 2023). To mitigate this, connectivity in $\mathcal{E}_{l+1}$ can be strengthened by adding $K$-hop edges.
@@
The upsampling process is represented by EXPAND, which is the inverse of the REDUCE function and transfers information from the coarser level back to the finer level. Most previous work generates coarser graphs either by using numerical software or by downsampling the input mesh through heuristic pooling strategies (Cao et al., 2023; Lino et al., 2022; Yu et al., 2024; Janny et al., 2023). This process is performed during the data preprocessing stage. The preprocessed hierarchy with the same input mesh topology is reused across different initial conditions and time steps.
Figure 1: The architecture of EvoMesh. Physical dynamics is modeled on multiple graph resolutions with adaptive structures, $\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{L}$, and are processed using their respective AMP layers. The DiffSELECT operation performs differentiable pooling to create coarser graphs with learnable downsampling probabilities. REDUCE and EXPAND integrate inter-level information using learned feature aggregation weights over the neighboring nodes. EvoMesh is trained end-to-end with one-step supervision.
@@ 3 Method
In this section, we introduce EvoMesh, a fully differentiable model that adaptively generates time-evolving graph hierarchies over the sequence, while simultaneously simulating the physical system over these learned hierarchical graphs. Figure 1 demonstrates an overview of the proposed model, which operates in an encode-process-decode pipeline. The encoder first maps the input field to a latent feature space $\mathbf{V}_{1}=\{\mathbf{v}_{i}|v_{i}\in\mathcal{V}_{1}\}$ at the original mesh resolution. Subsequently, we model the physical dynamics across the learned multi-scale graph hierarchies with adaptive graph structures.
In Section 3.1, we present the details of the AMP layer. In Section 3.2, we discuss the approach for learning context-aware graph hierarchies. In Section 3.3, we describe the inter-level downsampling and upsampling processes that incorporate AMP-based feature propagation. Finally, in Section 3.4, we outline the implementation details.
@@ 3.1 Anisotropic Message Passing
We introduce the AMP layer, which facilitates information propagation both within and between graph hierarchies, enabling EvoMesh to effectively capture local and long-range dependencies simultaneously.
As shown in Eq. (2), a common non-parametric aggregation in GNN-based mesh simulation is to use the summation for node update: $\mathbf{\hat{v}}_{i}=\phi^{v}\Big{(}\mathbf{v}_{i},\sum_{v_{j}\in\mathcal{N}_{v_{i}}}\mathbf{\hat{e}}_{ij}\Big{)}$, where $v_{j}\in\mathcal{N}_{v_{i}}$ denotes a neighboring node of $v_{i}$ in the graph.
To differentiate the contributions of neighboring nodes, the AMP layer employs learnable parameters $\phi^{w}$ to predict the anisotropic importance weight of edge feature $\mathbf{\hat{e}}_{ij}$ with respect to node $v_{i}$. These weights are then normalized across the neighborhood of $v_{i}$ using a softmax function:
$w_{ij}=\phi^{w}(\mathbf{e}_{ij},\mathbf{v}_{i},\mathbf{v}_{j}),\alpha_{ij}=\frac{\exp\left(w_{ij}\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(w_{ik}\right)}.$ (3)
@@
The proposed AMP layer enables the implicit assignment of varying contribution weights to the updated edge features within the same neighborhood. Analyzing the learned direction-specific weights in AMP further enhances interpretability. We adopt an MLP implementation for $\phi^{w}$, while alternative designs, such as graph attention (Veličković et al., 2018) or cross-attention (Vaswani et al., 2017), are also feasible. A detailed comparison is provided in Appendix C.2.
@@ 3.2 Differentiable Multi-Scale Graph Construction
In the DiffSELECT operation, we train the node update module $\phi^{v}$ based on anisotropic aggregated edge features to produce a 2-dimensional probability vector $\boldsymbol{\pi}_{i}^{l}$ for each node $v_{i}$. This vector $\boldsymbol{\pi}_{i}^{l}=(\pi_{i,0}^{l},\pi_{i,1}^{l})$ represents the probabilities of discarding or retaining node $v_{i}$ in the next-level coarser graph $\mathcal{G}_{l+1}$. We rewrite Eq. (4) as follows:
$\mathbf{\hat{v}}_{i}^{l},\ \boldsymbol{\pi}_{i}^{l}=\phi^{v}\left(\mathbf{v}_{i}^{l},\sum_{v_{j}\in\mathcal{N}_{v_{i}}}\alpha_{ij}^{l}\mathbf{\hat{e}}_{ij}^{l}\right).$ (5)
In the next step, we apply Gumbel-Softmax sampling (Jang et al., 2017) independently to each node, using the log-probabilities $(\log\pi_{i,0}^{l},\log\pi_{i,1}^{l})$ as logits. This produces a soft one-hot vector $\mathbf{z}_{i}^{l}=(z_{i,0}^{l},z_{i,1}^{l})$ for each node:
[…]
$=\frac{\exp\left((\log\pi_{i,k}^{l}+g_{i,k}^{l})/\tau\right)}{\sum_{k^{\prime}=0}^{1}\exp\left((\log\pi_{i,k^{\prime}}^{l}+g_{i,k^{\prime}}^{l})/\tau\right)},$
where $g_{i,k}^{l}$ is Gumbel noise sampled independently for each node, and $\tau$ is the temperature parameter controlling the smoothness of the sampling. In this way, the node set $\mathcal{V}_{l+1}$ is adaptively constructed based on node features from the finer graph level. The straight-through Gumbel-Softmax estimator provides a differentiable approximation to hard sampling, thereby facilitating end-to-end training. We implement the Gumbel-Softmax with temperature annealing to stabilize training, initially encouraging the exploration of hierarchies and gradually refining the selection process.
The edges $\mathcal{E}_{l+1}$ in the coarser graph $\mathcal{G}_{l+1}$ are constructed by connecting the selected nodes using the original graph’s edges $\mathcal{E}_{l}$. However, this process may result in disconnected partitions (see Appendix Figure 6). To address this issue, we enhance the connectivity in $\mathcal{E}_{l+1}$ by incorporating the $K$-hop edges during edge selection, defined as follows:
[…]
@@
Table 2: Quantitative comparison of the one-step and long-term prediction errors. We report the mean results over $3$ random seeds, with corresponding standard deviations detailed in Appendix C.7. Promotion denotes the improvement over the second-best model.
Model RMSE-1 (×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$)
Cylinder Airfoil Flag Plate Cylinder Airfoil Flag Plate
MGN (2021) 0.3046 77.38 0.3459 0.0579 59.78 2816 115.3 3.982
Lino et al. (2022) 3.9352 85.66 0.9993 0.0291 27.60 2080 118.2 2.090
BSMS-GNN (2023) 0.2263 71.69 0.5080 0.0632 16.98 2493 168.1 1.811
Eagle (2023) 0.1733 51.55 0.3805 0.0392 20.05 2344 127.7 7.797
HCMT (2024) 0.9190 48.62 0.4013 0.0295 23.59 3238 90.32 2.468
EvoMesh 0.1568 41.41 0.3049 0.0282 6.571 2002 76.16 1.296
Promotion 9.53% 14.8% 11.9% 3.10% 61.3% 3.75% 15.7% 28.5%